OpenStudy (anonymous):

what is the volume of a cylinder with a base area of 23cm^2 and a height of 5 cm? why answer is 115cm^3?

OpenStudy (dangerousjesse):

\(v=\pi r^2h\) \(a=\pi r^2\) Area of the base is \(23^2\) so we're left with \(v=23^2\times 5\)

OpenStudy (dangerousjesse):

Don't solve with the exponents- those stand for cubic centimeters.